# Observation of relaxation stages in a non-equilibrium closed quantum system: decaying turbulence in a trapped superfluid

### v1

@@ Abstract
In all our measurements, we keep the condensed fraction before excitation above $80\%$. For an excitation amplitude of $A=1.87\mu_{0}$, upon the repopulation of low momentum modes, it reaches a condensate fraction $\approx 20\%$. The introduced excitations eventually relax, resulting in considerable thermal excitations. The temperature of the condensate before the excitation is 50(6) nK, and the temperature of the cloud at $t\approx 370$ ms, when the condensate repopulation peak occurs, is 95(8) nK, such that the temperature increase is $\Delta T=45(10)$ nK. The conclusion is that the energy introduced by the excitation protocol for this amplitude still leaves the final cloud at a temperature that is less than the critical temperature $T_{c}\approx 130$ nK, which favors the condensate repopulation.
Figure S2: Illustration of the angular averaging procedure to compute the momentum distributions for the excitation amplitudes (a) $A=1.87\mu_{0}$ and (b) $2.30\mu_{0}$, in both cases for $t\approx 100$ ms. The dashed lines denote two $30^{\circ}$ angles centered around the major axis of the expanded cloud enclosing the region where $n(k)$ is calculated. (c),(d) The same information as (a) and (b), but expressed in polar coordinates. (e) Resulting angular averaged momentum distribution curves. The appearance of power-law behavior in the momentum range $5\ \mu\mathrm{m}^{-1}\leqslant k\leqslant 12\ \mu\mathrm{m}^{-1}$ supports the establishment of a turbulent state.
@@
Determining the universal exponents.— Although in the main text we focus on the NTFP present in the dynamics of the system subjected to an excitation amplitude of $1.87\mu_{0}$, we observe the universal scaling described by Eq. (1) in a total of eight excitation amplitudes in the interval $1.87\mu_{0}\leqslant A\leqslant 2.55\mu_{0}$.

### v2

@@ Acknowledgements.
Figure S2: Single shot illustrations of the angular averaging procedure to compute the momentum distributions for the excitation amplitudes (a) $A=1.87\mu_{0}$ and (b) $2.30\mu_{0}$, in both cases for $t\approx 100$ ms. The dashed lines denote two $30^{\circ}$ angles centered around the major axis of the expanded cloud enclosing the region where $n(k)$ is calculated. (c),(d) The same information as (a) and (b), but expressed in polar coordinates. (e) Resulting angular averaged momentum distribution curves. The appearance of power-law behavior in the momentum range $4\ \mu\mathrm{m}^{-1}\leqslant k\leqslant 12\ \mu\mathrm{m}^{-1}$ supports the establishment of a turbulent state.
Figure S3: Momentum distributions for an excitation amplitude of $A=1.87\mu_{0}$ during (a) the direct and (b) inverse particle cascades. These plots correspond to a larger momentum range than Figs. 4(b) and (e), where it is possible to see the particle transport from low to high momentum during the direct cascade and in the opposite direction for the inverse cascade.
Figure S4: Time evolution of $n(k\rightarrow 0,t)$ for the excitation amplitudes $A=1.70\mu_{0}$, $1.87\mu_{0}$, $2.04\mu_{0}$, $2.21\mu_{0}$, $2.30\mu_{0}$, and $2.55\mu_{0}$, panels (a)-(f), respectively.
@@
The direct cascade is characterized by particle transport from low- to high-momenta, while the inverse cascade occurs in the opposite direction. Both scenarios are illustrated in Fig. S3, where we present the time-evolving momentum distributions during both cascades. The values of $n(k\rightarrow 0,t)$ reported throughout the manuscript are the average of all values with $k\leqslant 0.33\ \mu\mathrm{m}^{-1}$. The pixel size in our absorption images corresponds to a momentum value of $0.066\ \mu\mathrm{m}^{-1}$.
The observation of the inverse cascade depends on a delicate balance. If too much energy is injected into the system, the condensate is quickly depleted and thermalized. On the other hand, if not enough energy is injected, the population of the high-momentum modes does not increase significantly, and identifying the inverse cascade is challenging.
Figure S4 shows the time evolution of $n(k\rightarrow 0,t)$ for the three excitation amplitudes depicted in Fig. 2 alongside three other excitation amplitudes, $A=2.04\mu_{0}$, $2.21\mu_{0}$, and $2.55\mu_{0}$. For $A=1.70\mu_{0}$, the excitation amplitude is so low that we do not observe characteristics of a turbulent system, such as a power-law behavior in the momentum distribution, nor compelling evidence of repopulation of the low-momentum modes. As the excitation amplitude progressively increases, the system shows evidence of condensate repopulation at $\approx 300$ ms. The most apparent evidence of this phenomenon was measured with an excitation amplitude of $1.87\mu_{0}$ at $t\approx 370$ ms, which is the one we thoroughly investigate in the manuscript. Using higher excitation amplitudes, where we also observed the repopulation of low momenta, the system fails to reach such high values of $n(k\to 0,t)$. For excitation amplitudes $A\geqslant 2.04\mu_{0}$, we also observed that when the repopulation occurs, it is always at times close to 300 ms, with a slight advancement relative to lower amplitudes, e.g., for $A=2.04\mu_{0}$ the peak occurs at $t\approx 300$ ms, while for $A=2.21\mu_{0}$ at $t\approx 260$ ms.
Determining the universal exponents related to the direct cascade.— Although in the main text we focus on the NTFP present in the dynamics of the system subjected to an excitation amplitude of $1.87\mu_{0}$, we observe the universal scaling described by Eq. (1) in a total of eight excitation amplitudes in the interval $1.87\mu_{0}\leqslant A\leqslant 2.55\mu_{0}$.
@@
Determining the universal exponents related to the inverse cascade.— We followed a similar procedure to the one described above, or in Ref. [28], to determine the universal exponents $\lambda$ and $\mu$. Compared to the procedure to extract $\alpha$ and $\beta$, a complication is that we must also determine the value of the blowup time $t_{b}$. We circumvent this problem by also varying $t_{b}$, and choosing the value which minimizes the uncertainties in $\lambda$ and $\mu$.
We start by defining the function
$\chi^{2}(\lambda,\mu)=\frac{1}{N_{t}^{2}}\sum_{t=t_{1}}^{t_{N_{t}}}\sum_{t_{\rm ref}=t_{1}}^{t_{N_{t}}}\chi^{2}_{\lambda,\mu}(t,t_{\rm ref}),$ (S6)
where we the sums run over all time instants $\{t_{1},\cdots,t_{N_{t}}\}$. The function $\chi^{2}_{\lambda,\mu}(t,t_{\rm ref})$ is given by
[…]
where $\sigma(k,t)$ corresponds to the standard deviation of the mean. The upper limit of the integration $k_{f}$ was varied, and we found that the exponents are insensitive to variations of $\delta k\approx 1.0\,\mathrm{\mu{m}^{-1}}$ around $k_{f}=6.0\,\mathrm{\mu{m}^{-1}}$.
Using the likelihood function
$L(\lambda,\mu)=\exp\left[-\frac{1}{2}\chi^{2}(\lambda,\mu)\right],$ (S8)
the exponents and their uncertainties are determined from a Gaussian fit of the marginal-likelihood functions, the same procedure used in Eqs. (S2) and (S3) with $\alpha$ and $\beta$ interchanged by $\lambda$ and $\mu$, respectively.
In Fig. S7(a), we show the likelihood function and its marginal likelihoods in the bottom and left panels for the case presented in the main text, a universal scaling window of $288$ ms $\leqslant t\leqslant 374$ ms. We obtained $\lambda=-1.5(5)$ and $\mu=-0.9(3)$ for a blowup time of $t_{b}=520(10)$ ms. If we consider a slightly larger temporal window, $266$ ms $\leqslant t\leqslant 374$ ms, we obtain the likelihoods illustrated in Fig. S7(b). The exponents are the same as the previous case but with more considerable uncertainties, $\lambda=-1.5(7)$ and $\mu=-0.9(4)$ for $t_{b}=560(10)$ ms. This indicates that the dynamical scaling region may be more restricted than indicated by only the $k\to 0$ component. The fit shown in Fig. 3 considers the functional form $n(k\to 0,t)\propto(t_{b}-t)^{\lambda}$ where we fixed the value of $\lambda$ at its theoretical estimate, ${\lambda}={\nu/(2-2\nu)}=-1.46$. We obtained $t_{b}=530(30)$ ms in accordance with the full scaling procedure considering finite values of the momentum.
In Fig. S8, we show the result of the dynamical scaling applied to the larger temporal window, $266$ ms $\leqslant t\leqslant 374$ ms, the equivalent of panel (e) and (f) of Fig. 4. Although it is possible to identify the collapse after the universal scaling procedure, the dispersion is more significant than the case we considered in the main text. This is in agreement with the larger uncertainties in the exponents if we increase the temporal window.
Figure S7: Likelihood functions for an excitation amplitude of $A=1.87\mu_{0}$ used to determine the exponents of the dynamical scaling related to the inverse cascade, Eq. (2). (a) The temporal window $288$ ms $\leqslant t\leqslant 374$ ms yields the exponents reported in the main text, $\lambda=-1.5(5)$ and $\mu=-0.9(3)$. (b) A slightly larger interval, $266$ ms $\leqslant t\leqslant 374$ ms, produces the same exponents with larger uncertainties, $\lambda=-1.5(7)$ and $\mu=-0.9(4)$.
Figure S8: (a) Momentum distributions for an excitation amplitude of $A=1.87\mu_{0}$ during $266$ ms $\leqslant t\leqslant 374$ ms, a larger interval than the one considered in the main text, Fig. 4(e). (b) Scaled momentum distributions using Eq. (2) with $\lambda=-1.5(7)$ and $\mu=-0.9(4)$.